While walking down the street, you stumble on a sphere of solid gold. The diameter of the sphere is 1.25 in. The density of gold is about 19.3 g/cm^3. If Gold sells for $850.00/oz, How much is the sphere worth? The equation for volume of a sphere is 3/4 3.14•radius^3, and there are 28.35g to an ounce.
1
Expert's answer
2021-03-02T05:59:28-0500

First we need to calculate the mass of golden sphere.

r=0.5d=0.625 in. = 1,5875 cm.

"v=\frac{4}{3}3.14r^{3}=\frac{4}{3}3.14\times1.5875^{3}=16.749 cm^{3}"

"m=v\rho=16.749\times19.3=323.2557.8g"

"cost/g = \frac{850}{28.35} = 29.98 /g"

price of golden sphere = "181.8\times29.98=9696.53 dollars"

Answer: 9696.5 dollars